degree of dissociation of weak monoprotic  acid can be given as 


(A) a= 1/[1+10(pKa+pH)]


(B) a=1/[1+10(pKa-pH)]


(C) a=1/[1+10(-pKa+pH)]


(D) a=1/[1+10(pKa/pH)]


please give detailed solution

Grade:12

1 Answers

AskiitianExpert Pramod-IIT-R
14 years ago
Dear Student MS. shefali, HA == H+ + A- c(1-a) ca ca Ka = ca2/(1-a) c= (1-a)/a2 * Ka...............(1) H+ = ca c = H+/a....................(2) by equation 1 and 2 H+/a == (1-a)/a2 * Ka H+ == (1-a)/a * Ka H+/Ka == (1-a)/a H+/Ka == (1/a)-1 (H+/Ka) +1 == (1/a) [10(-pH)]/[10(-pKa)] + 1 == (1/a) [10(pKa-pH) + 1] == (1/a) a=1/[1+10(pKa-pH)] So B option is correct. Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE & AIEEE preparation. All the best. Regards, Askiitians Experts Pramod Kumar IITR Alumni